OpenStudy (anonymous):

Medal/fan? :) Will you guys check my work? Also, I need help on two questions. Need Help: Given the binomials (x - 2), (x - 1), (x + 2), and (x - 4), which one is a factor of f(x) = x3 + 7x2 + 14x + 8? (x - 2) (x - 1) (x + 2) (x - 4) Given the binomials (x + 1), (x + 4), (x - 5), and (x - 2), which one is a factor of f(x) = 3x3 - 12x2 - 4x - 55? (x + 1) (x + 4) (x - 5) (x - 2)

OpenStudy (anonymous):

it would take some time ... to solve

OpenStudy (anonymous):

ok it is pretty simple .....

OpenStudy (anonymous):

Check My Work: What is the f of x over the g of x when f(x) = 6x3 - 19x2 + 16x - 4 and g(x) = x - 2? 6x2 - 7x + 2 (My answer) 3x2 - 9x + 8 6x2 - 7x + 2 - 8 over the quantity of x minus 2 3x2 - 9x + 8 - 8 over the quantity of x minus 2 What is the quotient when -3x3 + 5x + 14 is divided by x - 2? -3x2 - 6x - 7 - 3x2 - x + 12 (My answer) -3x2 + 6x - 7 + 28 over the quantity of x minus 2 -3x2 - x + 12 + 28 over the quantity of x minus 2 What is the quotient when x3 - 5x2 + 2x + 5 is divided by x - 2? x2 - 3x - 4 x2 - 7x + 16 x2 - 3x - 4 - 3 over the quantity of x minus 2 (my answer) x2 - 7x + 16 - 3 over the quantity of x minus 2

OpenStudy (anonymous):

why dont you just plug in the values of 2 ,1 , -2,-4 in the first question and see which one comes out as zero

OpenStudy (anonymous):

i think that is the simplest way to get it okay first you plug in 2

OpenStudy (anonymous):

you wont get it with 1 and 2 ... now for he negative values ..

OpenStudy (anonymous):

for the first one i would say x+2 do you know what is a factor ? it is something which when put to a equation you would get zero on the other side ...

OpenStudy (campbell_st):

just use the factor theorem if (x -2) is a factor then f(2) = 0 saves a lot of solving... it just requires substitution... whichever values, when substituted returns a zero, then that is a zero of the polynomial, and then you have a factor

OpenStudy (anonymous):

campbell .. you could help her right?

OpenStudy (campbell_st):

the 1st question is really straight forward if you know about the zeros of a polynomial... since every term in the cubic is positive, substituting a positive number will give a positive value... so start by selecting the binomial factor that would have a negative zero... and try that.

OpenStudy (campbell_st):

so here is the basics if (x -2) is a factor x = 2 is a zero substitute \[f(2) = (2)^3 + 7(2)^2 + 14(2) + 8\] its a positive answer... so not a factor... try (x + 2) which means x - 2 is a zero substitute and check \[f(-2) = (-2)^3 + 7(-2)^2 + 14(-2) + 8\] if you evaluate it you'll get f(-2) = 0 so (x + 2) is a binomial factor. hope this makes sense

OpenStudy (anonymous):

it does.

OpenStudy (campbell_st):

for your 2nd question (x+ 1) then if x = -1 is a zero show f(-1) = 0 (x +4) then x - 4 is a zero show f(-4) = 0 same process for the other binomial factors... as you can see, no difficult solving... hope this helps...

OpenStudy (anonymous):

What is the f of x over the g of x when f(x) = 6x3 - 19x2 + 16x - 4 and g(x) = x - 2? 6x2 - 7x + 2 (My answer) You did this correct (I just took the test)